OpenStudy (hannahc234):

Solve the quadratic equation 5x^2-8x=8-5x

OpenStudy (aveline):

How do we start?

OpenStudy (hannahc234):

We start by adding 5x to both sides

OpenStudy (aveline):

Right. We want one side to be equal to zero. So after we add 5x to both sides, we subtract 8 from both sides\[5x ^{2}-3x-8=0\]

OpenStudy (aveline):

What do we do next?

OpenStudy (hannahc234):

Factor?

OpenStudy (aveline):

Right

OpenStudy (caozeyuan):

why not just plug numbers into the solution equation, factoring seems to be kind of difficulty

OpenStudy (xcior12):

wassup u dome retrice c-u-n-t-s

OpenStudy (xcior12):

fck openstudy

OpenStudy (xcior12):

it sucks a ssss

OpenStudy (xcior12):

like ur mother

OpenStudy (aveline):

@mathmale We need a moderator

OpenStudy (xcior12):

so does this d-i-c-k

OpenStudy (aveline):

@caozeyuan Plugging numbers in may get you the right answer, but it would take forever. Not to mention that you might not even find all the possible answers.

OpenStudy (xcior12):

ave stop talking before buffy the white boy slayer come for ur retrice

OpenStudy (xcior12):

a\\ssss

OpenStudy (xcior12):

ill make yall shut the fck down going through all questions and messing with people

OpenStudy (welshfella):

and if you are studying quadratic equations you have to learn how to factor them.

OpenStudy (caozeyuan):

but the solution equation ganratees all solutions will be found, and factoring is hard espeicially if you have complex conjugate roots

OpenStudy (aveline):

@Hero @ganeshie8 We need a moderator here please

OpenStudy (xcior12):

oh u back did buffy the white boy slayer get ur a/////sss

OpenStudy (aveline):

@HannahC234 So sorry about this. Do you need help factoring?

OpenStudy (hannahc234):

Yes please

OpenStudy (aveline):

Ok, 5x^2-3x-8=0 For factoring a quadratic, we need to find what will add up to -3x and what will multiply to be negative 8: (5x-?)(x+?)

OpenStudy (hannahc234):

Okay. I'm still kinda lost

OpenStudy (aveline):

(5x-8)(x+1) is the next step. Do you know how to FOIL?

OpenStudy (hannahc234):

Yes

OpenStudy (aveline):

So if we FOIL, (5x-8)(x+1) equals: 5x^2+5x-8x-8 or 5x^2-3x-8 We know that (5x-8)(x+1) is equal to zero. So: 5x-8=0 AND x+1=0 Can you simplify these equations to find the values of x?

OpenStudy (hannahc234):

-1 and 8/5 would be the answer correct?

OpenStudy (mathmale):

Hannah: Substitute (-1) for x in the original equation. is the equation then true or false? That will tell you whether or not -1 is a solution.

OpenStudy (hannahc234):

So it would be -1?

OpenStudy (aveline):

"-1 and 8/5 would be the answer correct? " Both of those would are the right answers. If you plug them into the original equation, you'll see that they work.